OpenStudy (anonymous):

Solve the quadratic equation by factorisation 20x^2 - 31x - 9 = 0 No idea where to start, normally I would multiply the first by the last and then find the numbers that would multiply and give me 31. But having trouble!!!

OpenStudy (amistre64):

factor the product of 9(20) and see what subtract to get -31

OpenStudy (amistre64):

1,180; 2,90; 4,45 3,60; 6,30; 12,15 5, 36; ... 5 - 36 = -31

OpenStudy (amistre64):

(x-26/20) (x+5/20) ; simplify and if its still a fraction put the denom in front

OpenStudy (amistre64):

(x-36/20) (x+5/20) (x-9/5) (x+1/4) (5x-9) (4x+1)

OpenStudy (amistre64):

the 20 in the denom comes from the fact that we used it in the beginning to find our factors, so we have to divide it back out in the end

OpenStudy (amistre64):

but to get the zeros; just stop at the simplified fractions .. x = 9/5 and x = -1/4 are the zeros

OpenStudy (anonymous):

Thank you very much, just re-attempted it with your explanation (Y).